OpenStudy (anonymous):

What is the sum of a 17-term arithmetic sequence where the first term is 6 and the last term is -90?

OpenStudy (anonymous):

n(a1+an)/2

OpenStudy (anonymous):

Use the definition of an arithmetic sequence, and the required algebra to relate the first and last terms.

OpenStudy (anonymous):

idk how to do this at all

OpenStudy (anonymous):

what do i plug into what

OpenStudy (anonymous):

here..n=no.of terms. a1=first term. an=last term.

terenzreignz (terenzreignz):

First things first, @helpisuck What IS an arithmetic sequence (progression) ? U

terenzreignz (terenzreignz):

Or, on second thought, since you're just looking for the sum, anyway, @dlearner 's got it down... my bad :3

OpenStudy (anonymous):

so (17+6)/2?

OpenStudy (anonymous):

what do i do with the -90?

OpenStudy (nincompoop):

this website is a good high school prep review for the Regents, a NY mandate test for all HS to take before they graduate. http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm

OpenStudy (anonymous):

ima have to look at that later

OpenStudy (nincompoop):

you have to look at it now since the answers to the questions (except a direct answer to your problem) you have at the moment are in that link

terenzreignz (terenzreignz):

No, @helpisuck \[\Large \frac{n(a_1 + a_n)}{2}\] is the sum of the first n-terms of an arithmetic progression, true enough, but your 'plugging-in' is faulty... Now, let's replace the values one at a time, shall we? n here represents the number of terms you're adding, so what is n?

OpenStudy (anonymous):

i understand all that im just confused on how to plug it in. im not good with over reading on math i need it pretty much spelled out for me then explained why

OpenStudy (anonymous):

we'll make u spell it out :)

OpenStudy (anonymous):

say again lol

OpenStudy (anonymous):

basically

OpenStudy (anonymous):

i will have to but i cant now

OpenStudy (kenljw):

\[\sum_{k =1}^{n}k=\frac{ n(n + 1) }{ 2}\] negative the sum to 90 plus the sum to 6

OpenStudy (anonymous):

@helpisuck An arithmetic sequence is given in the form:\[\bf \left\{ a_n \right\} =\left\{ a,a+d,a+2d,a+3d...,a+(n-1)d \right\}\]We are given that the first term is 6 hence \(\bf a = 6\). We are also given the last term of this 17-TERM sequence as -90 hence \(\bf a+(17-1)d = 6+16d=-90\). Solving for 'd' gives us: \[\bf 6+16d=-90 \implies d = \frac{ -90-6 }{ 16 }=-6\]Now we have everything we need to compute the sum of the 17 terms of this sequence. The sum of any arithmetic sequence is given by:\[\bf S_n=\frac{ n }{ 2 }(2a+(n-1)d) \implies S_n=\frac{ n }{ 2 }(12+(n-1)(-6))\]Now simply evaluate \(\bf S_n\) when \(\bf n=17\). Can you do that?

OpenStudy (anonymous):

ive never seen the formula before either of them ive seen the first formlua that was posted but idk what to plug in where?

OpenStudy (anonymous):

can i ask you another question? if i have a 55 in my class and my final exam is coming up and it is worth like 320 points what would i need on the test to get a passing grade? (75)

OpenStudy (anonymous):

We are getting the sum of the first 17 terms. The formula for the first 'n' terms of an arithmetic sequence is given by \(\bf S_n=\frac{n}{2}(2a+(n-1)d)\) where \(\bf S_n\) represents the sum of the first \(\bf n\) terms. We figured from the given information that the first term is 6, i.e. \(\bf a=6\). We were also given that the last term, which is the 17th term is -90, i.e. \(\bf a+(17-1)d=6+16d=-90\). From here I solved for 'd' which gave me \(\bf d=-6\). Now we use the formula for \(\bf S_n\) which is the sum of the first n terms of an arithmetic sequence. Since we know what 'a' and 'd' are, we plug that in to our \(\bf S_n\) formula and we get:\[\bf S_n=\frac{ n }{ 2}(2(6)+(n-1)(-6))=\frac{ n }{ 2 }(12+(n-1)(-6))\]Now since we are finding the sum of the first 17 terms, we let \(\bf n = 16\) and evaluate \(\bf S_n\) accordingly. Can you do that now? @helpisuck

terenzreignz (terenzreignz):

Two things @helpisuck 1: We can't answer that ^ unless we know how 'heavy' exactly your final exam is... 2: Just do your best on that exam. Can't go wrong that way. If you can perfect it (I have faith in you ^_^) then perfect it... And that starts with understanding concepts ^ Don't worry if you think you're being spoonfed, we were all spoonfed before we learned to pick up our own cutlery :P

OpenStudy (anonymous):

@helpisuck First finish up this question and then make a new question..

OpenStudy (kenljw):

It depend on your performance prior to final and whether grades are on the standard bell curve, only so many A, B, C, D, F.

terenzreignz (terenzreignz):

@helpIsuck has left... Maybe it wasn't such a good idea to be so technical so early on @genius12 :3

OpenStudy (anonymous):

@terenzreignz Lol I was trying to make as much sense of the knowledge as I possibly could. I also assumed that he had some understanding of such sequences...